Ninety-eight feet of fencing is required to fence in a triangular ostrich pen. If the second largest side is three feet longer than the shortest side and the longest side is one foot shorter than twice the shortest side, what is the length of the largest side?
Answer by checkley77(12844)

second largest side=s+3
longest=2s-1
s+s+3+2s-1=98
4s=98-2
4s=962
s=96/4
s=24 feet for the short side.
second=24+3=27 feet
longest=2*24-1 feet
l=48-1
l=47
proof:
24+27+47=98
98=98